2.0 g of benzoic acid dissolved in 25.0 g of
benzene shows a depression in freezing point
equal to 1.62 K. What is the percentage
association of benzoic acid ? Kf for benzene is
4.9 K kg mol-1.
1
Expert's answer
2019-09-18T01:05:56-0400

1) Find the theoretical delta T

"\Delta T=K·\frac{m_{BA}·1000}{M_{BA}·m_{B}}=4.9·\frac{2·1000}{25·122.12}=3.21"

2)"i=\frac{\Delta T_{real}}{\Delta T}=\frac{1.62}{3.21}=0.5"

3)"i=1-\alpha (1+n)= 1-\alpha (1+0.5)" "\implies" "\alpha =\frac{1-i}{0.5}=\frac{0.5}{0.5}=1"

Answer: 100%